# On the class of coding optimality of human languages and the origins of Zipf’s law

### v4

@@ 1 Introduction
Nearly 7000 languages are spoken on Earth [1]. As languages from the same linguistic family resemble for sharing a close common ancestor, a better proxy for the diversity of languages is the number of families. That number, estimated at about 430, is still a large number [1]. In spite of such diversity, languages exhibit statistical patterns called linguistic laws. One of laws with strongest support is Zipf’s law of abbreviation, the tendency of more frequent words to be shorter [2], that is supported in 986 languages, namely 14% of the languages spoken on earth [3]. The strength of the law relies on the fact that it is found independently of the linguistic family, writing system (script), the unit (words or individual characters), modality (oral versus written), the unit of measurement (characters, syllables, strokes or duration) [4, 3, 5, 6, 7, 8]. Furthermore, this law is being found in a growing number of species with some interesting exceptions [9].
@@
It has been demonstrated that, if $l\right>$ is minimum, then the Kendall $\tau$ correlation between $p(i)$ and $l(i)$ cannot be positive, i.e. [10] $\tau(p(i),l(i))\leq 0$, which sheds light on the origins of the law of abbreviation.
Table 1: Distinct encodings of numbers from 1 to 6 with an alphabet formed by ${a,b}$.
@@
$1$ $aa$ $aa$ $b$
$2$ $ab$ $aa$ $aba$
$3$ $a$ $a$ $abb$
$4$ $b$ $b$ $aabaa$
$5$ $ba$ $ba$ $aabab$
$6$ $bb$ $bb$ $aabba$
In standard information theory, one assumes that $l(i)$ is the length of the code assigned to unit $i$. The code is a string of characters from some alphabet of size $N$. The problem of compression consists of finding the codes whose length minimizes $l\right>$. We use $l\right>_{min}$ to refer to the minimum value of $l\right>$. If no constraint is imposed on the possible codes, the minima of $l\right>$ are singular. If $l(i)=0$ is allowed, $l\right>$ is minimized by assigning the empty string to each unit. Then $l\right>_{min}=0$ and information transmission is impossible with empty strings. If $l(i)>0$ is required, $l\right>$ is minimized by assigning a string of just one character to each unit. Then $l\right>_{min}=1$ and $N$ has to be as large as the number of units for successful decoding, which, if possible, is eventually an overkill. For this reason, information theory is concerned about constraints on the mapping of units into codes, which are known as coding schemes [11]. In the non-singular coding scheme, every unit is assigned a distinct string. Suppose that $N=2$ and the alphabet is $\{a,b\}$. Table 1 shows a coding table that is non-singular and a coding table that is not non-singular because the string $aa$ is assigned to both unit 1 and unit 2. In the uniquely decodable coding scheme, a particular case of non-singular coding, the segmentation of a string resulting from the concatenation of codes must be unique given the coding table. For instance, $baba$ allows for many interpretations according to the non-singular encoding in Table 1: the segmentation $ba$, $ba$ implies that the units coded are $5$, $5$; the segmentation $b$, $a$, $b$, $a$ implies that the units coded are $4$, $3$, $4$, $3$, and so on. Therefore, the non-singular codes in Table 1 do not satisfy the uniquely decodable coding scheme. In contrast, the Elias gamma encoding [12] in Table 1 does so: the sequence $baba$ can only be segmented as $b$, $aba$ and then the units coded are certainly $1$, $2$.
Hereafter we assume $N>1$. Under optimal non-singular coding, $l\right>$ is minimized by [13, 10]
@@
while, under uniquely decodable encoding, $l\right>$ is minimized by [14, 11]
@@
for optimal uniquely decodable encoding. These soft versions have already been used in previous research on optimal coding with continuous magnitudes [15, 16] and can be justified as follows. First, Eq. 3 is well approximated by Eq. 5 as $\frac{N-1}{N}$ converges quickly to 1 as $N$ increases. Second, the solution of the constrained minimization of $l\right>$ by means of Lagrange multipliers produces directly Eq. 5 and Eq. 4 results a posteriori by imposing discreteness on $l(i)$ [11].
@@ 2 A new class of coding optimality
[…]
@@
$l\right>_{min}^{ud}=H=-\sum_{i}p(i)\log_{N}p(i).$ (9)
The following property indicates that, when Zipf’s law holds, the optimal code length under unique decodability is shifted linearly from the optimal code length under non-singular coding. By the same token, $H$ is shifted linearly from $\log_{N}i\right>$. The same applies to the coding efficiencies.
@@
Zipf’s rank-frequency law with exponent $\alpha$ and normalization factor $c$ is equivalent to a linear dependence between optimal code lengths with parameters $\alpha$ and $\beta=-\log_{N}c$ and $\alpha,\beta\geq 0$. First, $l_{min}^{ud}(i)$ and ł$\l_{min}^{ns}(i)$ are linearly dependent as
@@
[…]
@@
Dividing by $l\right>$ on the last result, one obtains Eq. 11. We have $\alpha\geq 0$ by definition of rank. To conclude that $\beta\geq 0$, one does not need to know the precise definition of $c$. Notice that $c=p(1)$ (Eq. 7). As $0<p(1)\leq 1$, $\beta>0$ follows. ∎
@@
(exponentiating with base $N$)
@@ 5.2 A new class of coding optimality
We presented the hypothesis that languages belong to a class of communication systems that are linearly displaced from these two kinds of optimal coding and thus must exhibit Zipf’s law. This is consistent with the view of Zipf’s law as a characteristic of the most compressed loss-less representations [38]. We have confirmed that the class contains at least three languages. The ubiquity of Zipf’s law in languages [2, 17] suggests this class would comprise many more languages. We cannot rule out the possibility that it does not cover all. A statistical analysis of a wider set of languages should be the target of future research.
@@ 5.3 The origins of Zipf’s law
where $\alpha$ and $\beta$ are Lagrange multipliers. If one assumes optimal non-singular coding (with $N\gg 1)$, then $l\right>$ becomes $\log_{N}i\right>$ and the solution to the maxent problem is Zipf’s law (Eq. 7) as already pointed out by Mandelbrot [13]. If one assumes $N=1$ or replaces $l\right>$ by the expected rank, that is simply, $i\right>$, it is well-known that the solution will be an exponential distribution [50]. Mandelbrot’s classic argument raises two concerns. The first one is the validity of the maxent argument for non-equilibrium systems such as language. One can argue that the maxent still provides the most unbiased distribution satisfying given constraints [50]. The second one is the validity of the assumption of optimal non-singular coding. Recall that $\eta^{ns}$ is on average $30\%$ [41]. Here we have put forward a class of languages that overcomes the limitations of Mandelbrot’s argument. The members of that class do not need to be coding optimally for Zipf’s law to surface, they only need to be linearly close to the optima. Modern optimization models about the origins of Zipf’s law have similar limitations [47, 38]. For instance, the proposal of a general relevance-resolution trade-off [38] (see Appendix A of [51] for a language-oriented summary of the framework) suggests, at first glance, an alternative general explanation to the origins of Zipf’s law that does not involve word lengths as ours. However, it assumes optimal uniquely decodable encoding when defining speaker and hearer costs and in particular when defining the cost of a word. Recall that $\eta^{ns}$ is on average $40\%$ [41]. In contrast, our proposal does not impose neither optimal coding nor a specific coding scheme. Instead, it uses optimal non-singular coding and optimal uniquely decodable encoding as reference points from which natural languages are hypothesized to deviate linearly from. While we aim to provide a unified explanation to linguistic laws via compression in communication systems, [38] has a much wider scope in terms of systems. We hope that our research stimulates a dialogue between these two complementary approaches.

### v5

@@ 1 Introduction
Nearly 7000 languages are spoken on Earth [1]. As languages from the same linguistic family resemble each other for sharing a close common ancestor, a better proxy for the diversity of languages is the number of families. That number, estimated at about 430, is still a large number [1]. In spite of such diversity, languages exhibit statistical patterns called linguistic laws. One of laws with strongest support is Zipf’s law of abbreviation, the tendency of more frequent words to be shorter [2], that is supported in 986 languages, namely 14% of the languages spoken on earth [3]. The strength of the law relies on the fact that it is found independently of the linguistic family, writing system (script), the unit (words or individual characters), modality (oral versus written), the unit of measurement (characters, syllables, strokes or duration) [4, 3, 5, 6, 7, 8]. Furthermore, this law is being found in a growing number of species with some interesting exceptions [9].
@@
It has been demonstrated that, if $\left<l\right$> is minimum, then the Kendall $\tau$ correlation between $p(i)$ and $l(i)$ cannot be positive, i.e. [10] $\tau(p(i),l(i))\leq 0$, which sheds light on the origins of the law of abbreviation.
Table 1: Distinct encodings of numbers from 1 to 6 with an alphabet formed by a,b{a,b}.
@@
11 aaaa aaaa $b$
22 abab aaaa abaaba
33 $a$ $a$ abbabb
44 $b$ $b$ aabaaaabaa
55 baba baba aababaabab
66 bbbb bbbb aabbaaabba
In standard information theory, one assumes that $l(i)$ is the length of the code assigned to unit $i$. The code is a string of characters from some alphabet of size $N$. The problem of compression consists of finding the codes whose length minimizes $\left<l\right$>. We use $\left<l\right>_{min}$ to refer to the minimum value of $\left<l\right$>. If no constraint is imposed on the possible codes, the minima of $\left<l\right$> are singular. If $l(i)=0$ is allowed, $\left<l\right$> is minimized by assigning the empty string to each unit. Then $\left<l\right>_{min}=0$ and information transmission is impossible with empty strings. If $l(i)>0$ is required, $\left<l\right$> is minimized by assigning a string of just one character to each unit. Then $\left<l\right>_{min}=1$ and $N$ has to be as large as the number of units for successful decoding, which, if possible, is eventually an overkill. For this reason, information theory is concerned about constraints on the mapping of units into codes, which are known as coding schemes [11]. In the non-singular coding scheme, every unit is assigned a distinct string. Suppose that $N=2$ and the alphabet is $\{a,b\}$. Table 1 shows a coding table that is non-singular and a coding table that is not non-singular because the string $aa$ is assigned to both unit 1 and unit 2. In the uniquely decodable coding scheme, a particular case of non-singular coding, the segmentation of a string resulting from the concatenation of codes must be unique given the coding table. For instance, babababa allows for many interpretations according to the non-singular encoding in Table 1: the segmentation $ba$, baba implies that the units coded are $5$, 55; the segmentation $b$, $a$, $b$, $a$ implies that the units coded are $4$, 33, 44, 33, and so on. Therefore, the non-singular codes in Table 1 do not satisfy the uniquely decodable coding scheme. In contrast, the Elias gamma encoding [12] in Table 1 does so: the sequence babababa can only be segmented as $b$, abaaba and then the units coded are certainly $1$, 22.
Hereafter we assume $N>1$. Under optimal non-singular coding, $\left<l\right$> is minimized by [13, 10]
@@ 44 $b$ $b$ aabaaaabaa
while, under uniquely decodable encoding, $\left<l\right$> is minimized by [14, 11]
@@
for optimal uniquely decodable encoding. These soft versions have already been used in previous research on optimal coding with continuous magnitudes [15, 16] and can be justified as follows. First, Eq. 3 is well approximated by Eq. 5 as $\frac{N-1}{N}$ converges quickly to 1 as $N$ increases. Second, the solution of the constrained minimization of $\left<l\right$> by means of Lagrange multipliers produces directly Eq. 5 and Eq. 4 results a posteriori by imposing discreteness on $l(i)$ [11].
@@ 2 A new class of coding optimality
$\eta=\frac{\left<l\right>_{min}}{\left<l\right>},$
measures the closeness to optimal non-singular coding. We use $\left<l\right>_{min}^{ns}$ and $\left<l\right>_{min}^{ud}$ to refer to the minimum value of $\left<l\right$> under non-singular ($ns$) coding and uniquely decodable ($ud$) encoding. Accordingly, we define $\eta^{ns}=\left<l\right>_{min}^{ns}/\left<l\right$> and $\eta^{ud}=\left<l\right>_{min}^{ud}/\left<l\right$>. Eq. 1 with $l(i)$ as in Eq. 5 yields the expected value of $\log_{N}i$, that is
$\left<l\right>_{min}^{ns}=\left<\log_{N}i\right>=\sum_{i}p(i)\log_{N}i.$ (8)
@@
$\left<l\right>_{min}^{ud}=H=-\sum_{i}p(i)\log_{N}p(i).$ (9)
The following property indicates that, when Zipf’s law holds, the optimal code length under unique decodability is shifted linearly from the optimal code length under non-singular coding. By the same token, $H$ is shifted linearly from $\left<\log_{N}i\right$>. The same applies to the coding efficiencies.
@@
Zipf’s rank-frequency law with exponent $\alpha$ and normalization factor $c$ is equivalent to a linear dependence between optimal code lengths with parameters $\alpha$ and $\beta=-\log_{N}c$ and $\alpha,\beta\geq 0$. First, $l_{min}^{ud}(i)$ and $\l _{min}^{ns}(i)$ are linearly dependent as
@@
Second, $\left<l\right>_{min}^{ud}=H$ and $\left<l\right>_{min}^{ns}=\left<\log_{N}i\right$> are also linearly dependent, i.e.
@@
Dividing both sides of the the last result by $\left<l\right$>, one obtains Eq. 11. We have $\alpha\geq 0$ by definition of rank. To conclude that $\beta\geq 0$, one does not need to know the precise definition of $c$. Notice that $c=p(1)$ (Eq. 7). As $0<p(1)\leq 1$, $\beta>0$ follows. ∎
@@
(exponentiating with base NN)
@@ 5.2 A new class of coding optimality
We presented the hypothesis that languages belong to a class of communication systems that are linearly displaced from these two kinds of optimal coding and thus must exhibit Zipf’s law. This is consistent with the view of Zipf’s law as a characteristic of the most compressed lossless representations [38]. We have confirmed that the class contains at least three languages. The ubiquity of Zipf’s law in languages [2, 17] suggests that this class would comprise many more languages. We cannot rule out the possibility that it does not cover all. A statistical analysis of a wider set of languages should be the target of future research.
@@ 5.3 The origins of Zipf’s law
where $\alpha$ and $\beta$ are Lagrange multipliers. If one assumes optimal non-singular coding (with $N\gg 1)$, then $\left<l\right$> becomes $\left<\log_{N}i\right$> and the solution to the maxent problem is Zipf’s law (Eq. 7) as already pointed out by Mandelbrot [13]. If one assumes $N=1$ or replaces $\left<l\right$> by the expected rank, that is simply, $\left<i\right$>, it is well-known that the solution will be an exponential distribution [50]. Mandelbrot’s classic argument raises two concerns. The first one is the validity of the maxent argument for non-equilibrium systems such as language. One can argue that the maxent still provides the most unbiased distribution satisfying given constraints [50]. The second one is the validity of the assumption of optimal non-singular coding. Recall that $\eta^{ns}$ is on average $30\%$ [41]. Here we have put forward a class of languages that overcomes the limitations of Mandelbrot’s argument. The members of that class do not need to be coding optimally for Zipf’s law to surface, they only need to be linearly close to the optima. Modern optimization models about the origins of Zipf’s law have similar limitations [47, 38]. For instance, the proposal of a general relevance-resolution trade-off [38] (see Appendix A of [51] for a language-oriented summary of the framework) suggests, at first glance, an alternative general explanation to the origins of Zipf’s law that does not involve word lengths as ours. However, it assumes optimal uniquely decodable encoding when defining speaker and hearer costs and in particular when defining the cost of a word. Recall that $\eta^{ns}$ is on average $40\%$ [41]. In contrast, our proposal does not impose either optimal coding or a specific coding scheme. Instead, it uses optimal non-singular coding and optimal uniquely decodable encoding as reference points from which natural languages are hypothesized to deviate linearly from. While we aim to provide a unified explanation to linguistic laws via compression in communication systems, [38] has a much wider scope in terms of systems. We hope that our research stimulates a dialogue between these two complementary approaches.